# Dropping Just a Handful of Preferences Can Change Top Large Language Model Rankings

### v1

@@ Abstract
We propose a method for evaluating the robustness of a widely used LLM ranking systemthe Bradley–Terry ranking systemto dropping a worst-case very small fraction of evaluation data. Our approach is computationally fast and easy to adopt. When we apply our method to matchups from two popular human-preference platforms, Chatbot Arena and MT-Bench, we find that the Bradley–Terry rankings of top-performing models are remarkably sensitive to the removal of a small fraction of evaluations. Our framework also identifies the specific evaluations most responsible for such ranking flips, allowing for inspections of these influential preferences. We observe that the rankings derived from MT-Bench preferences are notably more robust than those from Chatbot Arena, likely due to MT-bench’s use of expert annotators and carefully constructed prompts. Finally, we find that rankings based on crowdsourced human-evaluated systems are just as sensitive as those based on LLM-as-a-judge evaluations, where in both, dropping as little as $0.02\%$ of the total evaluations in the dataset can change the top-ranked model.
@@ 1 Introduction
Open evaluation platforms like Chatbot Arena have, in large part due to their openness, become a gold standard for assessing the capabilities of leading LLMs via human preference. These open platforms are now widely used by top LLM developers and companies to evaluate and design new models and benchmarks (Chiang et al., 2024a, ; Singh et al.,, 2025; Grattafiori et al.,, 2024; Hui et al.,, 2024; White et al.,, 2024).
@@
Such platforms rely on crowdsourced pairwise battles and human votes to compute model rankings (Lee et al.,, 2023; Bai et al.,, 2022). At the heart of these preference-based evaluation and alignment pipelines is the Bradley–Terry (BT) model (Bradley and Terry,, 1952; Sun et al.,, 2025), which is widely used to rank LLMs based on human feedback (Chiang et al., 2024a, ), train reward models for RLHF (Ouyang et al.,, 2022; Touvron et al.,, 2023; Xu et al.,, 2024), and route incoming queries to the most appropriate LLM or inference-time scaling strategy (Damani et al.,, 2024).
Recent controversies, however, have raised concerns about the trustworthiness of popular LLM evaluation platforms like Chatbot Arena (Chiang et al., 2024a, ). For instance, Min et al., (2025) demonstrate that Chatbot Arena is vulnerable to vote-rigging: by injecting just a few hundred manipulated votes (out of 1.7 million), attackers can significantly change the top model rankings. Singh et al., (2025) further identify issues such as data leakage and private testing practices that allow large, proprietary model developers to selectively report the best-performing versions of their models.
We study a slightly different type of untrustworthiness of LLM ranking systems in this work. That is: “Will the top rankings from LLM-evaluation platforms change upon dropping a very small fraction of the human (or AI) preference evaluations?” If it turns out that the answer is yes, this raises concerns about the generalizability and stability of the rankings produced by such systems. One might worry about whether these rankings, or the learned human preferences from these systems, actually generalize (Broderick et al.,, 2020).
This question motivates the need for a systematic way to assess the robustness of top rankings in BT-based evaluation systems to worst-case data dropping. However, no such method currently exists (beyond a brute-force combinatorial search over all possible small subsets of data) to test whether the model rankings on LLM evaluation systems are robust to the removal of a very small fraction of adversarially-chosen evaluations. [^1] This combinatorial search is computationally infeasible for large-scale platforms like Chatbot Arena.
@@
In order to avoid this computationally intractable combinatorial search, we turn to a recent line of works from statistics and theoretical computer science that design algorithms for assessing whether data analyses are robust to dropping a small, worst-case fraction of data points (Broderick et al.,, 2020; Kuschnig et al.,, 2021; Moitra and Rohatgi,, 2022; Freund and Hopkins,, 2023; Shiffman et al.,, 2023; Nguyen et al.,, 2024; Huang et al.,, 2024; Rubinstein and Hopkins,, 2025). One such method, the Approximate Maximum Influence Perturbation (AMIP), estimates how much a statistic of interest could change if a worst-case subset of the data were dropped (Broderick et al.,, 2020). We extend these ideas to develop a fast approximation method for assessing the robustness of rankings from LLM evaluation systems to worst-case data-dropping. We apply our method to two widely used sets of human preference matchups: Chatbot Arena and MT-Bench (Chiang et al., 2024a, ; Zheng et al.,, 2023). Although our methods are largely based on the ranking system used in Chatbot Arena and MT-Bench, there are a few differences (including tie-handling and post-processing) between the raw BT scores we analyze and the scores implemented on these platforms; see Section 3.1 for more details. We note that a previous work in computational biology (Shiffman et al.,, 2023) studies the sensitivity of gene set enrichment analyses to dropping just a few cells and that, importantly, this work also studies worst-case data-dropping robustness of rankings. However, while this previous work studies robustness in the ranking of p-values (specifically, p-values derived using the hypergeometric test), our work studies the robustness of rankings derived from the BT model (which can be thought of as elements in the coefficient vector of a logistic regression model).
@@
In Section 2, we formalize the setup for assessing worst-case data-dropping robustness in BT-based ranking systems, and in Section 3 we introduce a computationally efficient method for assessing this form of robustness in practice. In Section 4, we apply our robustness assessment method to investigate the robustness of LLM preference data.
Our code is publicly available at https://github.com/JennyHuang19/IsRankingRobust, including all scripts required to run our robustness assessing method and to reproduce the results presented in this paper.
@@ 2 Setup
We consider a preference-based ranking system akin to Chatbot Arena (Chiang et al., 2024a, ). There are in total $M$ language models. Any user can submit a prompt to be answered by a pair of language models, say models $i$ and $j$ for $i,j\in[M]$ with $i\neq j$. The user then determines if the response from model $i$ is better than that of model $j$. Suppose there are in total $N$ such comparisons; the data can be seen as a collection of tuples $\{(i_{n},j_{n},y_{n})\}_{n=1}^{N}$, with $y_{n}\in\{0,1\}$ being an indicator for whether in the $n$th match, model $i_{n}$ is preferred over model $j_{n}$. From a collection of preference data, the goal is to rank the language models.
Ranking with the Bradley–Terry model. The Bradley–Terry (BT) model is a classical statistical model used to rank players from binary preference data that is used by Chatbot Arena. In this model, each player (e.g., language model), $i$, is associated with a BT score, $\theta_{i}$, and the preferences are assumed to be generated as
$y_{n}\sim\text{Bernoulli}(\sigma(\theta_{i_{n}}-\theta_{j_{n}})),$ (1)
where the sigmoid function $\sigma(x)=1/(1+e^{-x})$. Note, since the “winning” probability depends on the difference between two players rather than on their raw scores, the scores are identified only up to a constant additive term. There are different ways to avoid this identifiability problem (Wu et al.,, 2022). We adopted the method of setting the first player’s score to be 0. We can cast this model as a logistic regression with a specially-structured design matrix. We denote the corresponding “design” vector of the $n$th comparison, $x_{n}\in\{-1,0,1\}^{M}$, a vector encoding which two players are being compared. That is, if the game is between players $i$ and $j$, then $x_{n}$ has a $1$ in the $i$th element, a 1-1 in the $j$th element, and 0 otherwise. Using this structure, we can rewrite the model as a logistic regression model with M1M-1 parameters corresponding to the scores of the players, $\bm{\theta}=(\theta_{1},\dots,\theta_{M})\in\mathbb{R}^{M}$ with $\theta_{1}=0$,
$y_{n}\sim\text{Bernoulli}(\sigma(x_{n}^{\top}\bm{\theta})).$ (2)
We fit the BT-model (i.e., estimate $\bm{\theta}$) by maximum likelihood,
$\hat{\bm{\theta}}$= $\operatorname*{arg\,min}_{\bm{\theta}:\theta_{1}=0}\;\sum_{n=1}^{N}\big{(}-y_{n}\log\sigma(x_{n}^{\top}\bm{\theta})$ (3)
$\qquad\qquad\quad-(1-y_{n})\log(1-\sigma(x_{n}^{\top}\bm{\theta}))\big{)}.$
@@
In the rest of this section, we turn to the goal of determining whether there exists a small fraction of data (e.g., matchups) that we can drop to change the ordering of the estimated BT scores.
Setup for Data-Dropping. We study whether dropping a small fraction $\alpha\in(0,1)$ (e.g., $\alpha=0.01$) of the preference data can change the ordering of the estimated BT scores. Broderick et al., (2020) define the Maximum Influence Perturbation as the largest possible change induced in a quantity of interest by removing at most $100\alpha\%$ of the data.
@@
$\hat{\bm{\theta}}(w)$= $\operatorname*{arg\,min}_{\bm{\theta}(w):\theta_{1}(w)=0}\;\sum_{n=1}^{N}w_{n}\big{(}-y_{n}\log\sigma(x_{n}^{\top}\bm{\theta})$ (4)
$\qquad\qquad\quad-(1-y_{n})\log(1-\sigma(x_{n}^{\top}\bm{\theta}))\big{)}.$
Setting $w=1_{N}$ (the all-ones vector) recovers the BT scores computed on the full data (e.g., the original arena), while setting $w_{n}=0$ corresponds to dropping the $n$th data point (e.g., a matchup). We define the set of all weight vectors corresponding to dropping at most an $\alpha$-fraction of the data as follows:
Definition 2.1 (Feasible Drop Set).
@@
Two-Player Arena. We begin by analyzing the robustness of an arena involving just two players (e.g., LLMs): player $i$ and player $j$. Without loss of generality, we assume [^2] If this assumption does not hold, the identities of $i$ and $j$ can be swapped. that player $i$ has the higher estimated BT score on the full data:
@@
Definition 2.2 (Top-1 Data-Dropping Robustness in Two-Player Arenas).
An arena consisting of players $i$ and $j$ is said to be top-1 robust at level $\alpha$ if there does not exist a data weighting $w\in W_{\alpha}$ such that the BT scores reverse under reweighting:
$\left\{w\in W_{\alpha}:\widehat{\theta}_{i}(w)<\widehat{\theta}_{j}(w)\right\}=\emptyset.$ (5)
To generalize this setup beyond a two-player arena, we introduce more notation.
$M$-Player Arena. We now extend this notion to arenas with $M$ players, for any $M\geq 2$. Let $\mathcal{T}(w):=\{\widehat{\theta}_{i}(w)\}_{i=1}^{M}$ denote the set of BT scores under weighting $w$.
Definition 2.3 (Top-$k$ Set).
@@
Definition 2.4 (Top-$k$ Data-Dropping Robustness in $M$-Player Arenas).
@@
Equation 7 is nontrivial to directly verify; to directly check, we have to test out dropping all possible small-fraction subsets of the arena, a combinatorial operation that is computationally intractable in practice.
In Section 3, we show that verifying whether Eq. 7 holds can be reduced to checking the robustness of a series of pairwise comparisons. Specifically, top-$k$ robustness as defined in Definition 2.4 can be checked by checking whether there exists a reweighting $w\in W_{\alpha}$ that flips the ranking of a pair $(i,j)$ such that $i$ is inside and $j$ is outside the top-$k$ set. We then can test if such flipping can happen by a continuous approximation of the discrete weights $w$. We detail this procedure in Section 3.

### v2

@@ Abstract
We propose a method for evaluating the robustness of widely used LLM ranking systems—variants of a Bradley–Terry model—to dropping a worst-case very small fraction of preference data. Our approach is computationally fast and easy to adopt. When we apply our method to matchups from popular LLM ranking platforms, including Chatbot Arena and derivatives, we find that the rankings of top-performing models can be remarkably sensitive to the removal of a small fraction of preferences; for instance, dropping just $0.003\%$ of human preferences can change the top-ranked model on Chatbot Arena. Our robustness check identifies the specific preferences most responsible for such ranking flips, allowing for inspection of these influential preferences. We observe that the rankings derived from MT-bench preferences are notably more robust than those from Chatbot Arena, likely due to MT-bench’s use of expert annotators and carefully constructed prompts. Finally, we find that neither rankings based on crowdsourced human evaluations nor those based on LLM-as-a-judge preferences are systematically more sensitive than the other.
@@ 1 Introduction
Open evaluation platforms like Chatbot Arena (Chiang et al., 2024b) have, in large part due to their openness, become a gold standard for assessing the capabilities of leading LLMs via human preference. These open platforms are now widely used by top LLM developers and companies (The Wall Street Journal, 2024; Bloomberg, 2025) to evaluate and design new models and benchmarks (Grattafiori et al., 2024; Hui et al., 2024). Such platforms rely on crowdsourced pairwise battles and human votes to compute model rankings (Lee et al., 2023; Bai et al., 2022).
Figure 1: Our method (i) tests whether AI leaderboard rankings remain stable upon dropping small fractions of data and (ii) pinpoints the specific data points (e.g., preferences) that drive ranking flips.
@@
At the heart of these preference-based evaluation pipelines is the Bradley–Terry (BT) model (Bradley and Terry, 1952), which is widely used to rank LLMs based on human feedback (Chiang et al., 2024b). The BT model is also used to train reward models for RLHF (Ouyang et al., 2022; Touvron et al., 2023; Xu et al., 2024; Sun et al., 2025) and route queries to the most appropriate LLM or inference-time scaling strategy (Damani et al., 2025).
A growing body of work has called into question the trustworthiness of LLM leaderboards, showing that they are vulnerable to adversarial attacks: a few hundred injected votes can change top rankings on Chatbot Arena (Min et al., 2025), attackers can identify model outputs to systematically upvote or downvote targets (Huang et al., 2025b), LLM-judges can be easily gamed (Zheng et al., 2025; Raina et al., 2024), and issues such as data leakage or selective reporting further undermine leaderboard reliability (Singh et al., 2025).
In this work, we study a different type of untrustworthiness of LLM ranking systems. That is: “Will the top rankings from LLM-evaluation platforms change upon dropping a very small fraction of the human (or AI) preference evaluations?” A positive answer would raise concerns about the stability and generalizability of rankings produced by such systems. Our notion of non-robustness differs from those of Min et al. (2025); Huang et al. (2025b); Zhao et al. (2025) in two major respects. First, it occurs at a different place in the process, at the data analysis step after data has been collected (including from malicious or apathetic users). Second, it does not require adversarial intent. Our notion is more concerned with statistical robustness, namely of a ranking learned from data to dropping a small fraction of the data. While we do aim to find a worst-case fraction, the intent is to provide an upper bound on the degree of non-robustness.
@@
Our question posed above motivates the need for a systematic way to assess the robustness of top rankings in BT-based evaluation systems to worst-case data dropping. However, no such method currently exists, beyond a brute-force combinatorial search over all possible small subsets of data. [^1] This combinatorial search is computationally infeasible for large-scale platforms like Chatbot Arena. In order to avoid this computationally intractable search, we turn to a recent line of works from statistics and theoretical computer science that design algorithms for assessing whether data analyses are robust to dropping a small, worst-case fraction of data points (Broderick et al., 2020; Kuschnig et al., 2021; Moitra and Rohatgi, 2023; Freund and Hopkins, 2023; Shiffman et al., 2023; Nguyen et al., 2024; Huang et al., 2025a; Rubinstein and Hopkins, 2025). One such method, the Approximate Maximum Influence Perturbation (AMIP), estimates how much a statistic of interest could change if a worst-case subset of the data were dropped (Broderick et al., 2020). We extend these ideas to develop a fast approximation method for assessing the robustness of rankings from LLM evaluation systems to worst-case data-dropping.
We apply our method to assess several popular LLM ranking platforms, including Chatbot Arena and derivatives (Chiang et al., 2024b; Zheng et al., 2023; Miroyan et al., 2025; Vichare et al., 2025; Chou et al., 2025) and find most to be non-robust to dropping a very small fraction of votes.
@@
In Section 2, we formalize the setup for assessing worst-case data-dropping robustness in BT-based ranking systems, and in Section 3 we introduce a computationally efficient method for assessing this form of robustness in practice (Figure 1). In Section 4, we apply our robustness assessment method to investigate the robustness of several LLM leaderboards.
@@ 2 Setup
We consider a preference-based ranking system akin to Chatbot Arena (Chiang et al., 2024b). There are in total $M$ language models. Any user can submit a prompt to be answered by a pair of language models. Let the $n$th such prompt be sent to models $i_{n}$ and $j_{n}$ for $i_{n},j_{n}\in[M]:=\{1,\ldots,M\}$ with $i_{n}\neq j_{n}$. The user then determines if the response from model $i_{n}$ is better than that of model $j_{n}$, or is tied. Suppose there are in total $N$ such comparisons; the $n$th comparison can be seen as a tuple $(i_{n},j_{n},y_{n})$, with $y_{n}\in\{W,L,T\}$ for whether in the $n$th match, model $i_{n}$ is preferred over model $j_{n}$ (a win, $W$), $j_{n}$ is preferred over $i_{n}$ (a loss, $L$), or the two models are similar (a tie, $T$). From a collection of preference data, the goal is to rank the language models.
Ranking with the (unweighted) Bradley–Terry model. The Bradley–Terry (BT) model is a classical statistical model used to rank players from binary match outcomes when there are only wins and losses, $y_{n}\in\{W,L\}$. In this model, each player (e.g., language model), $i$, is associated with a BT score, $\theta_{i}$, and the outcomes are modeled as
$I_{y_{n}=W}\sim\text{Bernoulli}(\sigma(\theta_{i_{n}}-\theta_{j_{n}})),$ (1)
where the sigmoid function $\sigma(x)=1/(1+e^{-x})$ and $I$ is the indicator function. Note, since the “winning” probability depends on the difference between two players’ scores rather than on their raw scores, the scores are identified only up to a constant additive term. There are different ways to avoid this identifiability problem (Wu et al., 2022). Chatbot Arena chooses to set mixtral-8x7b-instruct-v0.1 as the reference model, assigning it a fixed score of 1,1141{,}114. Chatbot Arena computes the BT-scores (i.e., the estimates of $\bm{\theta}=(\theta_{1},\dots,\theta_{M})$) for the unweighted BT-model by maximum likelihood,
$\hat{\bm{\theta}}$= $\operatorname*{arg\,max}_{\bm{\theta}:\theta_{1}=0}\;\sum_{n=1}^{N}\big(I_{y_{n}=W}\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+I_{y_{n}=L}\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}}))\big).$ (2)
Ranking with the weighted Bradley–Terry model to handle ties. The classic BT model cannot handle ties. To handle ties, Chatbot Arena adds weights to Equation 2, counting a tie as both a win and a loss (Chiang et al., 2024b).222“Chatbot Arena Leaderboard Calculation (Bradley–Terry model)” Colab notebook:https://colab.research.google.com/drive/1KdwokPjirkTmpO_P1WByFNFiqxWQquwH. In the weighted BT model, one specifies a weight for wins and losses, $w_{WL}$, and a weight for ties $w_{T}$. That is, we estimate BT scores by maximizing the weighted likelihood,
$\hat{\bm{\theta}}$= $\operatorname*{arg\,max}_{\bm{\theta}:\theta_{1}=0}\;\sum_{n=1}^{N}\big[w_{WL}I_{y_{n}=W}\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+w_{WL}I_{y_{n}=L}\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}}))$ (3)
$\qquad\qquad\quad+w_{T}I_{y_{n}=T}\big(\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}}))\big)\big].$
As done on Chatbot Arena, we use $w_{WL}=2$ and $w_{T}=1$. This choice can be interpreted as each win or loss counting as two matches of the same outcome, and a tie counting as one win and one loss. They also suggested an alternative treatment of dropping all ties and using the unweighted BT model, which corresponds to $w_{WL}=1$ and $w_{T}=0$.
@@
Postprocessing in Chatbot Arena. Chatbot Arena applies a linear transformation to the learned BT scores (Chiang et al., 2024a). They use $\texttt{SCALE}=400$, $\texttt{INIT\_RATING}=1{,}000$, and a further shift ANCHOR_SHIFT to produce the displayed scores:
$\texttt{ELO}_{i}=\texttt{SCALE}\cdot\widehat{\theta}_{i}+\texttt{INIT\_RATING}+\texttt{ANCHOR\_SHIFT}.$
The final constant (ANCHOR_SHIFT) shifts all the $\texttt{ELO}_{i}$ scores so that a specific reference model has a certain score. Chatbot Arena uses mixtral-8x7b-instruct-v0.1 as the reference model, assigning it a fixed score of 1,1141{,}114. We use the same reference model in our analysis of Chatbot Arena; however, we assign the model a fixed score of 0 (a design choice that does not impact rankings). We note that the affine transformation does not affect model rankings since it is strictly monotonic and does not affect our proposed procedure since linear transformations can commute with first-order Taylor expansion.
Setup for Data-Dropping. We study whether dropping a small fraction $\alpha\in(0,1)$ (e.g., $\alpha=0.01$) of the preference data can change the ordering of the estimated BT scores. Broderick et al. (2020) define the Maximum Influence Perturbation as the largest possible change induced in a quantity of interest by removing at most $100\alpha\%$ of the data.
@@
$\hat{\bm{\theta}}(w)$= $\operatorname*{arg\,max}_{\bm{\theta}:\theta_{1}=0}\;\sum_{n=1}^{N}w_{n}\big[w_{WL}I_{y_{n}=W}\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+w_{WL}I_{y_{n}=L}\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}}))$ (4)
$\quad\quad\qquad\qquad\quad+w_{T}I_{y_{n}=T}\big(\log\sigma(\theta_{i_{n}}-\theta_{j_{n}})+\log(1-\sigma(\theta_{i_{n}}-\theta_{j_{n}}))\big)\big].$
Setting $w=1_{N}$ (the all-ones vector) recovers the BT scores computed on the full data (e.g., the original arena), while setting $w_{n}=0$ corresponds to dropping the $n$th data point (e.g., a matchup). We define the set of all weight vectors corresponding to dropping at most an $\alpha$-fraction of the data as follows.
Definition 1 (Feasible Drop Set).
@@
We begin by analyzing the robustness of the ordering of BT scores between a pair of players, $i$ and $j$. Without loss of generality, we assume [^3] If this assumption does not hold, the identities of $i$ and $j$ can be swapped. that player $i$ has the higher estimated BT score on the full data:
@@
We now extend this notion to an arena with $M$ players, for any $M\geq 2$. Let $\mathcal{T}(w):=\{\widehat{\theta}_{i}(w)\}_{i=1}^{M}$ denote the set of BT scores under weighting $w$.
Definition 2 (Top-$k$ Set).
@@
Definition 3 (Top-$k$ Data-Dropping Robustness).
@@
Notice that Equation 6 is nontrivial to directly verify; to check directly, we have to test out dropping all possible small-fraction subsets of the arena, a combinatorial operation that is computationally intractable in practice.
In Section 3, we show that verifying whether Equation 6 holds can be reduced to checking the robustness of a series of pairwise comparisons. Specifically, top-$k$ robustness as defined in Definition 3 can be checked by checking whether there exists a reweighting $w\in W_{\alpha}$ that flips the ranking of a pair $(i,j)$ such that $i$ is inside and $j$ is outside the top-$k$ set. We then can test if such flipping can happen by using a continuous approximation of the discrete weights $w$ (also known as “approximate data-dropping”) to identify a promising candidate subset of influential preferences, dropping these, recomputing the BT-based rankings, and observing whether the rankings change. We detail this procedure in Section 3.